Frozen (quarter-pound)hamburger patties cost $5.80 for 12. How many can be purchased for $7.00?
Show your work
Thanks from Precious

Found 3 solutions by fractalier, jmg, vihits13:
Answer by fractalier(2101)

Set up a proportion...
$5.80/12 = $7.00/x
now cross multiply and solve
5.8x = 84
x = 84/5.8
x = 14.48
but since no one will sell you .48 of a patty...the practical answer is...
14

Answer by jmg(22)

The first thing you need to do is find out how must one patty would cost.
So you divide 5.80 by 12.
This gives you $0.48.
It asks for how many you could buy for $7.00
So divide $7.00 by $0.48
This gives you 14.58, well that tells you that could buy 14 hamburger patties.
you can check with a calculator. If you round up and buy 15, you spend too much money.
Hope this helps!

Answer by vihits13(20)

You can solve this problem by setting up an proportion. 12 over $5.80 should be equal to the number of patties we can purchase with $7.00 over $7.00.
We will give the number of patties we can purchase with $7.00 a value of x.
12/5.80 = x/7 This is the proportion. Next we crosst-multiply.
12*7 = 5.80*x We now have an equation.
84 = 5.80x This is the simplified version.
14.48 = x Divided by 5.80 to get x.
It says that x equals 14.48 but you can't have part of a pattie left over so we can buy aprroximately 14 patties with money left over with $7.00.